# High Dimensional Distributed Gradient Descent with Arbitrary Number of Byzantine Attackers

### v2

@@ 15       Update parameter $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta g(\mathbf{w}_{t})$;
The above results show that, as long as the number of worker machines $m$ grows proportionally with $d$, and $N_{A}$ grows slightly with $d$, then the learning error does not increase with dimensionality. Hence, compared with previous works [15, 17, 18], our new method is more suitable in high dimensional problems.
A drawback of our approach and the corresponding theoretical analysis is that we have not derived a desirable bound under sparse setting, which means that $d\gg m$, i.e. the dimensionality is much larger than the number of worker machines. In this case, theoretical value of $\lambda_{c}$ in (27) can be very large. This problem also happens in [19] for the case with less than half worker machines attacked. However, despite such drawback in sparse settings, in our numerical experiments in section 6, we have shown that the practical performance is still better than existing approaches. Nevertheless, it is hopeful to explore robust aggregators in sparse settings [35, 44, 42].
6 Numerical Results
@@
1.Master only. This means that we only use the gradient values from the auxiliary clean data stored in the master:
$g_{MasterOnly}(\mathbf{w})=\frac{1}{N_{A}}\sum_{j=1}^{N_{A}}\nabla f(\mathbf{w},\mathbf{Z}_{ij}).$ (35)
We use this method as a baseline, in order to show the benefits of combining clean samples with the untrusted gradient information from worker machines.
2.Distance based filtering. This method comes from [15], which has two different versions depending on whether $q$ is known. Here we just assume that $q$ is known, in order to compare our method with a better baseline.
The idea is that among all $m$ gradient vectors from worker machines, pick $m-q$ closest one, and then calculate the weighted average:
$g_{BasicFiltering}(\mathbf{w})=\frac{N_{A}g_{MasterOnly}+n\sum_{i\in\mathcal{N}(g_{MasterOnly}(\mathbf{w}),m-q)}\mathbf{Y}_{i}}{N_{A}+n(m-q)},$ (36)
in which $\mathcal{N}(g_{MasterOnly}(\mathbf{w}),m-q)$ means the $m-q$ nearest neighbors of $g_{MasterOnly}$ among corrupted gradient vectors from worker machines, $\{\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m}\}$.
3.Zeno. This method was proposed in [17], which filter gradients by stochastic descent scores, and was improved in [18]. [18] introduces several new modifications. Here we use the first order expansion of stochastic descent score mentioned in [18]. Other new techniques in [18] are not used, since they aimed at the suitability in asynchronous scenarios, which is out of scope of our work. In particular, this method assigns a score
$Score(\mathbf{Y}_{i},g_{MasterOnly})=\gamma\langle g_{MasterOnly},\mathbf{Y}_{i}\rangle-\rho\|\mathbf{Y}_{i}\|_{2}^{2},$ (37)
and then use the average gradients from $m-q$ worker machines with highest scores:
$g_{Zeno}(\mathbf{w})=\frac{1}{m-q}\mathbf{Y}_{(i)},$ (38)
in which $\mathbf{Y}_{(i)}$ is the gradient vector with $i$-th highest score according to (37).
@@
We would like to remark here that our method relies on auxiliary clean dataset, thus we only compare with previous methods that also use auxiliary samples. Other methods designed for $q<m/2$, such as Krum [12], geometric median-means [13], coordinate-wise median or trimmed mean [14], and recent high dimensional methods [19], are not shown here due to unfair comparison.
6.1 Synthesized Data
@@
in which $\mathbf{U}_{i},\mathbf{w}^{*}\in\mathbb{R}^{d}$, $\mathbf{w}^{*}$ is the true parameter vector, $W_{i}$ is the i.i.d noise following normal distribution $\mathcal{N}(0,1)$. Now we randomly generate $\mathbf{w}^{*}$ with each coordinate follows normal distribution $\mathcal{N}(0,2)$. We generate $N=50,000$ samples. For each sample, all components of $\mathbf{U}_{i}$ are i.i.d following $\mathcal{N}(0,1)$, and $V_{i}$ are calculated using (39). These samples are distributed into $m=500$ worker machines. Moreover, we have $N_{A}=50$ auxiliary clean samples. The results with $q=350$ and $q=150$ Byzantine machines are shown in Fig.1 and Fig.2, respectively. In both experiments, we show the results with different dimensionality $d=25,50,75,100,150,200$. Here we use a simple attack strategy. If a worker is attacked, each component of the gradient vector follows $\mathcal{N}(0,0.2)$.
(a) $d=25$.
(b) $d=50$.
(c) $d=75$.
(d) $d=100$.
(e) $d=150$.
(f) $d=200$.
Figure 1: Experiment results with synthesized data under linear model, with $q/m=0.7$.
From Fig.1, with $70\%$ worker machines attacked, if $d=25$, then the convergence of our method (red solid curve) is nearly the same as previous methods including distance based filtering and Zeno. The advantage of our method becomes clear under with $d=50$. From (d)-(f), if $d\geq 100$, then the weight vector does not converge reasonably using previous methods. However, our new approach is still effective.
(a) $d=25$.
(b) $d=50$.
(c) $d=75$.
(d) $d=100$.
(e) $d=150$.
(f) $d=200$.
Figure 2: Experiment results with synthesized data under linear model, with $q/m=0.3$.
Now we move on to the case with only $30\%$ worker machines being attacked. The results are shown in Fig.2. Unlike the case with $q>m/2$, here distance based filtering, Zeno and our method are all effective. However, our method is still better than previous methods.
6.2 Real data
Here we use MNIST dataset [45] to test the performance of robust gradient aggregators. MNIST dataset has $60,000$ images for training and $12,000$ images for testing. The size of each image is $28\times 28$.
The model is a neural network with one hidden layer between input and output. The size of hidden layer is $32$. In each experiment, we first randomly select $N_{A}=50$ images samples as the auxiliary clean dataset. The remaining samples are distributed into $m=500$ worker machines evenly. The gradients are obtained by backpropagating cross entropy loss function. The results for $q=350$ and $q=150$ are shown in Fig.3 (a) and (b), respectively.
(a) $q/m=0.7$.
(b) $q/m=0.3$.
@@ 6.2 Real data
In (a), majority of worker machines are attacked. In this case, distance based filtering can not capture correct gradient information in first $10$ iterations. The accuracy then begins to improve, but the performance is not competitive in general. Zeno performs much better. Our understanding is that the gradient descent score (37) actually selects gradient vectors whose directions are close to the gradient from auxiliary data. In neural networks with complex loss landscape, this rule may be better than filtering based entirely on distances. However, both methods still suffer from curse of dimensionality. The red solid line in the figure shows that our method performs significantly better than these previous methods. (b) shows the performances with only $30\%$ workers attacked. The results show that our method still performs better than previous methods, although not as obvious as the situation with $q/m>1/2$.
Previous results [15, 17, 18] use a much larger auxiliary dataset in their experiments. However, we only use $N_{A}=50$. As a result, the accuracy scores of existing methods in our experiments are worse than these early results in general. With large $d$ and small $N_{A}$, auxiliary dataset is not enough to train a model. In these case, traditional filtering techniques can improve the performance but the effects are limited. However, our new method performs significantly better.
7 Conclusion
@@
This paper solves the problem of high dimensional distributed learning problem under arbitrary number of Byzantine attackers. We primarily focus on the case $q>m/2$, but our method also apply in the case with $q<m/2$. To begin with, we have proposed a new method for semi-verified mean estimation, which combines a small clean dataset and a large corrupted dataset to estimate the statistical mean. We have also conducted theoretical analysis under both additive and strong contamination model. The results show that the new method is minimax rate optimal. We have then applied the semi-verified mean estimation method into the aggregator function in distributed learning. Compared with existing methods, the performance of our method is nearly the same as existing approaches under low dimensionality. In high dimension problems, our method is significantly better. Numerical results validate our theoretical analysis.
Appendix A Proof of Theorem 1
@@
([28], Proposition B.1) With probability at least $1-e^{-\alpha N/64}$, there exists a set $G_{0}\subset S_{0}^{*}$ with $|G_{0}|\geq\alpha N/2$, and
$\left\|\frac{1}{|G_{0}|}\sum_{i\in G_{0}}(\mathbf{Y}_{i}-\mu^{*})(\mathbf{Y}_{i}-\mu^{*})^{T}\right\|_{op}\leq 8\sigma^{2}\left(1+\frac{2d}{\alpha N}\right).$ (40)
With this lemma, we can pick $G_{0}$ with $|G_{0}|\geq\alpha N/2$ and satisfy (40). $G_{0}$ can be intuitively understood as ’good data’, which means that samples in $G_{0}$ are neither corrupted nor outliers in clean samples. Note that with a small probability $e^{-\alpha N/64}$, this is impossible. In this case, just let $G_{0}$ be a set of $\lceil\alpha N/2\rceil$ samples randomly selected from $S_{0}^{*}$.
@@
$\hat{\mu}(A)=\mathbf{P}\mu(A)+(\mathbf{I}-\mathbf{P})\mu(S),$ (41)
@@
$\|\hat{\mu}-\mu^{*}\|_{2}^{2}\leq 3\|\mathbf{P}(\mu(A)-\mu^{*})\|_{2}^{2}+3\|(\mathbf{I}-\mathbf{P})(\mu(S)-\mu(G))\|_{2}^{2}+3\|(\mathbf{I}-\mathbf{P})(\mu(G)-\mu^{*})\|_{2}^{2}.$ (42)
@@
Bound of $\|\mathbf{P}(\mu(A)-\mu^{*})\|_{2}^{2}$. We prove the following lemma:
@@
$\mathbb{E}[\|\mathbf{P}(\mu(A)-\mu^{*})\|_{2}^{2}]\leq\frac{\sigma^{2}p}{N_{A}}.$ (43)
@@
$\mathbb{E}[\|\mathbf{P}(\mu(A)-\mu^{*})\|_{2}^{2}]$ $=$ $\mathbb{E}[\operatorname{tr}(\mathbf{P}(\mu(A)-\mu^{*})(\mu(A)-\mu^{*})^{T}\mathbf{P})]$ (44)
$\leq$ $\frac{1}{|A|}\mathbb{E}[\operatorname{tr}(\sigma^{2}\mathbf{P})]$
@@
in which $\operatorname{tr}$ denotes trace of a matrix. For the last step, recall that $N_{A}=|A|$ is the size of auxiliary dataset, and $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}$, with $\mathbf{U}_{p}$ being the matrix of first $p$ column (step 1 in Algorithm 1).
@@
Under the following conditions: (1) (40) holds; (2) there exists an $\epsilon\in(0,1)$, such that
$\lambda_{c}$$\geq$ $32\sigma^{2}\left(1+\frac{2d}{\alpha N}\right)\frac{1+\epsilon}{1-\epsilon},$ (45)
@@
(3) In all iterations in Algorithm 1, denote $L_{t}\subset S_{t}$ as removed samples in step $t$, $G_{t}=G_{0}\cap S_{t}$ is the set of remaining good data after $t$ iterations, $L_{Gt}=L_{t}\cap G_{t}$, then for all iterations,
@@
The proof is shown in Appendix A.1. ∎
@@
$\text{P}\left(\exists t,\frac{|L_{Gt}|}{|G_{t}|}>\frac{1+\epsilon}{|G_{t}|}\sum_{i\in G}\frac{\tau_{i}}{\tau_{\max}}\text{ or }\frac{|L_{t}|}{|S_{t}|}<\frac{1-\epsilon}{|S_{t}|}\sum_{i\in S}\frac{\tau_{i}}{\tau_{\max}}\right)\leq 4N\exp\left[-\frac{1}{16}\lambda_{c}m\alpha^{2}N^{\frac{1}{3}}\epsilon^{2}\right].$ (50)
@@
The proof is shown in section A.3. ∎
If at least one of conditions (40), (47), (48) are violated, then due to the prefilter step (step 1) in Algorithm 1, the maximum distance between any two samples is at most $2N^{1/3}$. Therefore, $\|(\mathbf{I}-\mathbf{P})(\mu(S)-\mu(G))\|_{2}^{2}\leq 4N^{2/3}$ always hold. With Lemma 1 and 4, (64) is violated with exponential decaying probability. As a result,
$\mathbb{E}[\|(\mathbf{I}-\mathbf{P})(\mu(S)-\mu(G))\|_{2}^{2}]\leq\frac{2\lambda_{c}}{\alpha}+4N^{\frac{2}{3}}\left[e^{-\frac{1}{64}\alpha N}+4N\exp\left(-\frac{1}{16}\lambda_{c}m\alpha^{2}N^{\frac{1}{3}}\epsilon^{2}\right)\right].$ (51)
@@
The proof is shown in Appendix A.4. ∎
@@
$\mathbb{E}[\|(\mathbf{I}-\mathbf{P})(\mu(G)-\mu^{*})\|_{2}^{2}]=\frac{\lambda_{c}}{2\alpha}+4e^{-\frac{1}{64}\alpha N}N^{\frac{2}{3}}.$ (53)

### v3

@@ Introduction
The above results show that, as long as the number of worker machines $m$ grows proportionally with $d$, and $N_{A}$ grows slightly with $d$, then the learning error does not increase with dimensionality. Hence, compared with (cao2019distributed; xie2019zeno; xie2020zeno++), our new method removes the $\sqrt{d}$ dependence and is thus more suitable to high dimensional problems.
Numerical Results
@@
1. Master only. This means that we only use the gradient values from the auxiliary clean data stored in the master:
$g_{Master}(\mathbf{w})=\frac{1}{N_{A}}\sum_{j=1}^{N_{A}}\nabla f(\mathbf{w},\mathbf{Z}_{ij}).$ (33)
This method is used as a baseline, in order to show the benefits of combining clean samples with the untrusted gradient information from worker machines.
2. Distance based filtering. This method comes from (cao2019distributed). Here we just assume that $q$ is known. Among all $m$ gradient vectors from worker machines, this method picks mqm-q closest one, and then calculate the weighted average:
$g_{DBF}(\mathbf{w})=\frac{N_{A}g_{Master}+n\underset{i\in\mathcal{N}_{m-q}(g_{Master}(\mathbf{w}))}{\sum}\mathbf{Y}_{i}}{N_{A}+n(m-q)},$ (34)
in which $\mathcal{N}_{m-q}(g_{Master}(\mathbf{w}))$ means the mqm-q nearest neighbors of $g_{Master}$ among corrupted gradient vectors from worker machines, $\{\mathbf{Y}_{1},\ldots,\mathbf{Y}_{m}\}$.
3. Zeno. This method was proposed in (xie2019zeno). After that, an improved version was provided in (xie2020zeno++), called Zeno++. Here we use the first order expansion of stochastic descent score mentioned in (xie2020zeno++). In particular, this method assigns a score
$Score(\mathbf{Y}_{i},g_{Master})=\gamma\langle g_{Master},\mathbf{Y}_{i}\rangle-\rho\|\mathbf{Y}_{i}\|_{2}^{2},$ (35)
and then use the average gradients from mqm-q worker machines with highest scores: $g_{Zeno}(\mathbf{w})=(1/(m-q))\sum_{i=1}^{m-q}\mathbf{Y}_{(i)}$, in which $\mathbf{Y}_{(i)}$ is the gradient vector with $i$-th highest score according to (35).
@@
For all these four methods, we implement two types of attack: random attack, in which $\mathbf{Y}_{i}\sim\mathcal{N}(0,\sigma_{attack}^{2}\mathbf{I})$, and sign-flip attack, which flips the sign of original gradient vector, such that $\mathbf{Y}_{i}=-\mathbf{X}_{i}$.
We would like to remark here that our method relies on auxiliary clean dataset, thus we only compare with previous methods that also use auxiliary samples. Other methods designed for $q<m/2$, such as Krum (blanchard2017machine), geometric median-means (chen2017distributed), coordinate-wise median or trimmed mean (yin2018byzantine), and recent high dimensional methods (zhu2023byzantine), are not shown here due to unfair comparison.
(a) Random attack, $d=20$.
(b) Sign-flip attack, $d=20$.
(c) Random attack, $d=50$.
(d) Sign-flip attack, $d=50$.
(e) Random attack, $d=100$.
(f) Sign-flip attack, $d=100$.
Figure 2: Experiment results with synthesized data under linear model, with $q/m=0.8$.
Synthesized Data
@@
in which $\mathbf{U}_{i},\mathbf{w}^{*}\in\mathbb{R}^{d}$, $\mathbf{w}^{*}$ is the true parameter vector, $W_{i}$ is the i.i.d noise following normal distribution $\mathcal{N}(0,1)$. Now we randomly generate $\mathbf{w}^{*}$ with each coordinate following normal distribution $\mathcal{N}(0,2)$. We generate $N=50,000$ samples. For each sample, all components of $\mathbf{U}_{i}$ are i.i.d following $\mathcal{N}(0,1)$, and $V_{i}$ are calculated using (36). These samples are distributed into $m=500$ worker machines. The learning rate is set to be $\eta=5\times 10^{-3}$. Moreover, there are $N_{A}=50$ auxiliary clean samples. The results with $q=400$ (i.e. $\alpha=0.2$) and $q=100$ (i.e. $\alpha=0.8$) Byzantine machines are shown in Figure 2 and Figure 3, respectively. We set $p=5$ for $q=400$, and $p=2$ for $q=100$. In both experiments, we show the results with different dimensionality $d=20,50,100$. We run experiments with both random attack and sign-flip attack, in which $\sigma_{attack}=1$ for the former one.
(a) Random attack, $d=20$.
(b) Sign-flip attack, $d=20$.
(c) Random attack, $d=50$.
(d) Sign-flip attack, $d=50$.
(e) Random attack, $d=100$.
(f) Sign-flip attack, $d=100$.
Figure 3: Experiment results with synthesized data under linear model, with $q/m=0.2$.
(a) Random attack, $q/m=0.8$.
(b) Sign-flip attack, $q/m=0.8$.
(c) Random attack, $q/m=0.2$.
(d) Sign-flip attack, $q/m=0.2$.
@@
Figure 2 (a) and (b) show that with $80\%$ Byzantine workers, if $d=20$, then the loss curves of our method (green solid curves in the figure) are nearly the same as previous methods, including distance based filtering and Zeno. With $d=50$, our method begins to outperform existing methods, which is shown in (c) and (d). If the dimensionality is further increased to $d=100$, then the advantage of our method becomes clearer. This result agrees well with our theoretical analysis.
Now we move on to the case with only $20\%$ worker machines being attacked. The results are shown in Figure 3. The result shows that even in the case with less than half machines being attacked, our method still performs comparable to or better than existing approaches.
Real data
Here we use MNIST dataset (lecun1998mnist) to test the performance of robust gradient aggregators. In MNIST dataset, there are $60,000$ training images and $12,000$ testing images, with each image has a size of $28\times 28$.
The model is a neural network with one hidden layer between input and output. The size of hidden layer is $32$. In each experiment, we first randomly select $N_{A}=50$ samples as the auxiliary clean dataset. The remaining samples are distributed into $m=500$ worker machines evenly. The gradients are obtained by backpropagating cross entropy loss function. The results for $q=400$ (i.e. $\alpha=0.2$) and $q=100$ (i.e. $\alpha=0.8)$ under both random and sign-flip attack are shown in Figure 4, respectively. For random attack, $\sigma_{attack}=0.01$.
From Figure 4 (a)-(d), our new method outperforms existing approaches in general, for both random attack and sign-flip attack, especially with $q/m=0.8$. For the case with $q/m=0.2$, our method is slightly better than others, but the advantage becomes less obvious. This result is natural since our design is primarily for the case with more than half workers being Byzantine. Furthermore, we would like to remark that Zeno appears to perform better than distance-based filtering. Our understanding is that the gradient descent score (35) actually selects gradient vectors whose directions are close to the gradient from auxiliary data. In neural networks with complex loss landscape, this rule may be better than filtering based entirely on distances.
Conclusion
@@
This paper solves the problem of high dimensional distributed learning problem under arbitrary number of Byzantine attackers. Firstly, we have proposed a new method for semi-verified mean estimation, which combines a small clean dataset and a large corrupted dataset to estimate the statistical mean. We have also conducted theoretical analysis under both additive and strong contamination model. The results show that the new method is minimax rate optimal. We have then applied the semi-verified mean estimation method into the aggregator function in distributed learning. Compared with existing methods, the performance of our method is nearly the same as existing approaches for low dimensional problems. With higher dimensionality, our method performs significantly better. Finally, numerical results validate the effectiveness of our new method.
Acknowledgments
This work is supported by National Natural Science Foundation of China under Grants (62072106), the Open Research Projects of the Key Laboratory of Blockchain Technology and Data Security of the Ministry of Industry and Information Technology for the year 2025 (Grant KT20250015), the Natural Science Foundation of China (General Program, No.62571529), the Science and Technology Innovation Platform Project of Fujian Province, China (Grant 2023-P-003), and the Industry-University Cooperative Project for Colleges and Universities of Fujian Province, China (Grant 2024H6027).
Appendix A Appendix
In the appendix, we first present a review of related work. Next, we compare our direct semi-verified mean estimation approach with the indirect one based on list-decodable mean estimation. Finally, we provide formal proofs of the theorems.
Related Work
Robust mean estimation
Distributed learning relies on accurate estimation of the statistical mean of gradients gathered from worker machines. Therefore, we provide a brief review of robust mean estimation first. Estimation with corrupted samples has been an important task in statistics since 1960s (huber1964robust; huber2004robust). Traditional robust estimation methods such as trimmed mean behaves badly as dimensionality increases. An exception is Tukey median (tukey1975mathematics), which is sample efficient in high dimensional spaces. However, Tukey median is not practical, since the computation can not finish in polynomial time. (lai2016agnostic; diakonikolas2016robust; diakonikolas2017being) proposed convex programming and filtering methods for high dimensional robust mean estimation. Since then, a flurry of works has emerged, focusing on improving the time complexity (cheng2019high), extending to sparse problems (balakrishnan2017computationally; diakonikolas2020outlier), and improving the performance under stronger tail assumption with sum of squares method (hopkins2018mixture). These methods work if the proportion of contaminated samples is less than $1/2$.
If majority of samples are corrupted, then it is impossible to estimate the mean with a single solution. Suppose that the fraction of clean samples is ensured to be at least $\alpha$, i.e. $q/m<1-\alpha$, then the adversary can produce $\Omega(1/\alpha)$ clusters of points, with one of them composed of clean samples, while others are corrupted. In this case, there is no way to determine which cluster is correct. Therefore, in recent works, the goal of mean estimation is relaxed to generating a list of hypotheses, with the guarantee that at least one of the hypotheses is close to the ground truth $\mu^{*}$. This paradigm is called list-decodable learning. Typically, the length of generated hypotheses list need to be controlled within $O(1/\alpha)$. We hope to make the minimum estimation error among the hypotheses list to be as low as possible. (charikar2017learning) proposed an SDP-based method, achieving error of $O(\sigma\sqrt{\ln(1/\alpha)/\alpha})$. For spherical Gaussians, a multi-filter method can achieve a better rate (diakonikolas2018list). Several recent works improves the time complexity of list-decodable mean estimation. (diakonikolas2021list) proposed SIFT (Subspace Isotropic Filtering), which achieves nearly PCA time complexity. (cherapanamjeri2020list) introduced an optimization based approach with linear time complexity.
Apart from the untrusted data, if we also have some auxiliary clean samples, then it is possible to determine one estimated value from the hypotheses list, which is close to the ground truth with high probability. (charikar2017learning) named it semi-verified learning, and claims that with $K$ hypotheses, one needs $O(\ln K)$ clean samples to determine the correct one.
Federated Learning with less than half Byzantine machines
Suppose that there are $m$ worker machines, among which $q$ of them being attacked. Each worker machine has $n$ samples. The total size of dataset is $N=mn$. Here we briefly review the previous works with $q<m/2$.
(blanchard2017machine) proposed Krum, which collects gradient vectors from these $m$ worker machines, and then select a vector that is closest to its mqm-q neighbors. It is expected that the selected one is close to the ground truth. However, the error bound is suboptimal. (chen2017distributed) uses a geometric median of mean method, with a suboptimal error rate of $\sqrt{qd/N}$. (yin2018byzantine) made the first step towards optimal statistical rates of Byzantine robust federated learning. Two methods are proposed: coordinate-wise median and coordinate-wise trimmed mean. Under the assumption that the gradient has bounded skewness, the coordinate-wise median method achieves $\tilde{O}(q\sqrt{d}/(m\sqrt{n})+d/\sqrt{nm}+\sqrt{d}/n)$. If $n\gtrsim m$, then this method is optimal in $q$, $n$ and $m$. Under the subexponential tail assumption, the coordinate-wise trimmed mean achieves $\tilde{O}(d(q/(m\sqrt{n})+1/\sqrt{nm}))$, which is also optimal. However, these two methods are not optimal in $d$. As a result, in large scale models, the performance of these models are not desirable.
Recently, (zhu2023byzantine) shows that by using several recent high dimensional robust mean estimation methods (diakonikolas2016robust; diakonikolas2017being; diakonikolas2020outlier; steinhardt2018robust; zhu2022a; zhu2022b; zhu2022c; diakonikolas2023algorithmic) as the gradient aggregator functions, the error rate of federated learning becomes nearly optimal not only in $m$, $n$, $q$, but also in dimensionality $d$, up to a logarithm factor.
Federated learning with arbitrary number of Byzantine machines
The condition that less than half worker machines are attacked may not hold in practice. In this case, the methods mentioned above no longer works. If $q>m/2$, without additional information, it is impossible to provide a precise solution. (cao2019distributed) provides a simple method with the help of a clean auxiliary dataset. The basic idea is to calculate the gradient vector only using the auxiliary dataset as a baseline, and then filter out gradient vectors uploaded from worker machines that are too far away from such baseline. A slightly different method was proposed in (regatti2020bygars), which introduces reputation scores as the weights for gradient. However, the error rate of these methods is $\|\hat{\mathbf{w}}-\mathbf{w}^{*}\|=O(\sqrt{d/N_{A}})$, with $N_{A}$ being the size of auxiliary dataset. This rate is not significantly improved compared with learning with auxiliary data only, and is undesirable in high dimension problems. A significantly different method called Zeno was proposed in (xie2019zeno), which formulates a stochastic descent score to filter out adversarial workers. (xie2020zeno++) proposed Zeno++, an improvement to Zeno. The main advantage of Zeno++ is the suitability under asynchronous setting. Moreover, by computing a first order approximation of the gradient descent score, this method significantly improves the computational efficiency. However, the error bound is also significantly suboptimal. In particular, the error rate of Zeno and Zeno++ is still $O(\sqrt{d/N_{A}})$ (this comes from calculating the term $V_{3}$ in Theorem 1 in (xie2020zeno++)), which also scales badly with the increase of dimensionality. To the best of our knowledge, our work is the first attempt to solve Byzantine tolerant distributed learning problem in high dimensions.
Comparison with Indirect Method Based on List-decodable Mean Estimation
The SIFT method proposed in (diakonikolas2021list) is an efficient approach for list-decodable mean estimation. Given the ratio of trusted samples $\alpha$, SIFT outputs a list of $O(1/\alpha)$ hypotheses whose minimum distance to $\mu^{*}$ is $O(1/\alpha)$. From (charikar2017learning), with $O(\ln(1/\alpha))$ verified samples, the list-decodable method can be converted to semi-verified method.
The main difference of our method with the indirect approach mentioned above is that after obtaining the projection matrix, SIFT projects a list of random samples, i.e. calculating $\mathbf{P}\mathbf{X}_{0}+(\mathbf{I}-\mathbf{P})\mathbf{X}_{i}$ as one of hypothesis, in which $i$ is randomly selected. However, we only calculate one output, i.e. $\hat{\mu}=\mathbf{P}\mathbf{X}_{0}+(\mathbf{I}-\mathbf{P})\mu(S)$. There are also some differences in details of matrix calculation.
Now we compare our direct method with the indirect approach mentioned above. For a fair comparison, we let $n=1$, then under additive contamination model, (17) becomes
$\mathbb{E}[\left\lVert\hat{\mu}-\mu^{*}\right\rVert_{2}]\lesssim\sigma\alpha^{-\frac{1}{2}}\left(\frac{1}{\sqrt{N_{A}}}+1\right)\sim\sigma\alpha^{-\frac{1}{2}},$ (37)
as long as $1/\alpha\lesssim p\lesssim N_{A}/\alpha$.
Our method has several advantages. Firstly, according to (charikar2017learning), the indirect method using SIFT requires $O(\ln(1/\alpha))$ auxiliary clean samples, while we only need at least one clean sample. Secondly, the parameter selection is easier, since SIFT requires $p\sim 1/\alpha$, while our method only requires $1/\alpha\lesssim p\lesssim N_{A}/\alpha$, which is a wider constraint. In practice, $\alpha$ is usually unknown and setting an appropriate $p$ is challenging. Our direct method partially alleviates this issue.
Proof of Theorem 1
@@ Appendix A Appendix
((charikar2017learning), Proposition B.1) With probability at least $1-e^{-\alpha m/64}$, there exists a set $G_{0}\subset S_{0}^{*}$ with $|G_{0}|\geq\alpha m/2$, and
$\left\|\frac{1}{|G_{0}|}\sum_{i\in G_{0}}(\mathbf{Y}_{i}-\mu^{*})(\mathbf{Y}_{i}-\mu^{*})^{T}\right\|_{op}\leq 8\sigma^{2}\left(1+\frac{2d}{\alpha m}\right).$ (38)
With this lemma, we can pick $G_{0}$ with $|G_{0}|\geq\alpha m/2$ and satisfy (38). $G_{0}$ can be intuitively understood as ’good data’, which means that samples in $G_{0}$ are neither corrupted nor outliers in clean samples. Note that with a small probability $e^{-\alpha m/64}$, this is impossible. In this case, just let $G_{0}$ be a set of $\lceil\alpha m/2\rceil$ samples randomly selected from $S_{0}^{*}$.
@@
$\hat{\mu}(A)=\mathbf{P}\mathbf{X}_{0}+(\mathbf{I}-\mathbf{P})\mu(S),$ (39)
@@
$\|\hat{\mu}-\mu^{*}\|_{2}^{2}\leq 3\|\mathbf{P}(\mathbf{X}_{0}-\mu^{*})\|_{2}^{2}+3\|(\mathbf{I}-\mathbf{P})(\mu(S)-\mu(G))\|_{2}^{2}+3\|(\mathbf{I}-\mathbf{P})(\mu(G)-\mu^{*})\|_{2}^{2}.$ (40)
@@
Bound of $\|\mathbf{P}(\mathbf{X}_{0}-\mu^{*})\|_{2}^{2}$. We prove the following lemma:
@@
$\mathbb{E}[\|\mathbf{P}(\mathbf{X}_{0}-\mu^{*})\|_{2}^{2}]\leq\frac{\sigma^{2}p}{N_{A}}.$ (41)
Proof.
@@
$\mathbb{E}[\|\mathbf{P}(\mathbf{X}_{0}-\mu^{*})\|_{2}^{2}]$ $=$ $\mathbb{E}[\operatorname{tr}(\mathbf{P}(\mathbf{X}_{0}-\mu^{*})(\mathbf{X}_{0}-\mu^{*})^{T}\mathbf{P})]$ (42)
$\leq$ $\frac{1}{N_{A}}\mathbb{E}[\operatorname{tr}(\sigma^{2}\mathbf{P})]$
@@
in which $\operatorname{tr}$ denotes trace of a matrix. For the last step, recall that $\mathbf{P}=\mathbf{U}_{p}\mathbf{U}_{p}^{T}$, in which $\mathbf{U}_{p}\in\mathbb{R}^{d\times p}$ is the first $p$ columns of the matrix $\mathbf{U}$, whose columns are eigenvectors of $V(S)$. Hence $\operatorname{tr}(\mathbf{P})=p$. ∎
@@
Under the following conditions:
(a) (38) holds;
(b) there exists an $\epsilon\in(0,1)$, such that
$\lambda_{c}$ $\geq$ $32\frac{\sigma^{2}}{n}\left(1+\frac{2d}{\alpha m}\right)\frac{1+\epsilon}{1-\epsilon},$ (43)
@@
(c) In all iterations in Algorithm 2, denote $L_{t}\subset S_{t}$ as removed samples in step $t$, $G_{t}=G_{0}\cap S_{t}$ is the set of remaining good data after $t$ iterations, $L_{Gt}=L_{t}\cap G_{t}$, then for all iterations,
@@
The proof is shown in Appendix A. ∎
@@
$\text{P}\left(\exists t,\frac{|L_{Gt}|}{|G_{t}|}>\frac{1+\epsilon}{|G_{t}|}\sum_{i\in G}\frac{\tau_{i}}{\tau_{\max}}\text{ or }\frac{|L_{t}|}{|S_{t}|}<\frac{1-\epsilon}{|S_{t}|}\sum_{i\in S}\frac{\tau_{i}}{\tau_{\max}}\right)\leq 4m\exp\left[-\frac{1}{8}\alpha\lambda_{c}m^{\frac{1}{3}}\epsilon^{2}\right].$ (48)
@@
The proof is shown in section A. ∎
If at least one of conditions (38), (45), (46) are violated, then due to the prefilter step in Algorithm 2, the maximum distance between any two samples is at most $2m^{1/3}$. Therefore, $\|(\mathbf{I}-\mathbf{P})(\mu(S)-\mu(G))\|_{2}^{2}\leq 4m^{2/3}$ always hold. With Lemma 1 and 4, (62) is violated with exponential decaying probability. As a result,
$\mathbb{E}[\|(\mathbf{I}-\mathbf{P})(\mu(S)-\mu(G))\|_{2}^{2}]\leq\frac{2\lambda_{c}}{\alpha}+4m^{\frac{2}{3}}\left[e^{-\frac{1}{64}\alpha m}+4m\exp\left(-\frac{1}{8}\alpha\lambda_{c}m^{\frac{1}{3}}\epsilon^{2}\right)\right].$ (49)
@@
The proof is shown in Appendix A. ∎
@@
$\mathbb{E}[\|(\mathbf{I}-\mathbf{P})(\mu(G)-\mu^{*})\|_{2}^{2}]=\frac{\lambda_{c}}{2\alpha}+4e^{-\frac{1}{64}\alpha m}m^{\frac{2}{3}}.$ (51)